Given C = 150 + 0.50Y; I = 200 - 400i; Mt = 0.25Y; Ma = 50-100i and Ms = 180 for the money and goods markets. (Ms = Money Supply, Ma = Request for Money for speculative motives, Mt = Request for Money for transaction motives)
 
If the MPC (Marginal Propensity to Consume) rises to 0.6 and I equals 140, how much is the change in the money supply needed to maintain an investment of 140?
 
a. 32.52.
b. 36.25.
 c. 216.25.
 d. 200.50.
 e. 53.76.
 
the answer is B why?
Answer #1

Provided in the question,

Y = C + I

Y = 150 + 0.5Y + 200 - 400i

Here, 0.5 is the MPC

Y - 0.5Y = 350 - 400i

Y = 700 - 800i

This is the IS side of the equation.

Now, the LM side can be seen as,

Ms = Ma + Mt

Ms = 0.25Y + 50 - 100i

180 = 0.25Y + 50 - 100i

Now, put the equation of Y in this,

180 = 0.25(700 - 800i) + 50 - 100i

180 = 175 - 200i + 50 - 100i

300i = 175 + 50 - 180

i = 45/300

i = 0.15

Now, I is given as I = 200 - 400i, keep interest rate constant to keep Investment constant.

Changing MPC to 0.6

Y = 150 + 0.6Y + 200 - 400i

Y - 0.6Y = 350 - 400i

Y = 875 - 1000i

Put this in LM,

Ms = 0.25(875 - 1000i) + 50 - 100i

Ms = 218.75 - 250i + 50 - 100i

Ms = 268.75 - 350i

Keeping the interest rate constant to keep investment constant,

Ms = 268.75 - 350(0.15)

Ms = 268.75 - 52.5

Ms = 216.25

Initial Ms was 180 and new Ms is 216.25. The difference would be 216.25 - 180 = 36.25

Therefore, the answer is B.